Q1. Suppose Random variable X is # of jobs students applied with minimum 0 to maximum 23 that are uniformly distributed.

  1. Find P(2
  1. Find P(x=2)
  1. Find P(x>22)
Answer #1

Solution :

Given that,

a = 0

b =23

(B)P(x=2)=0.0000

probability=0.0000

(C)P(x > c) = (b - c) / (b - a)

P(x > 22) = (23 - 22) / (23 - 0)=0.0435

probability=0.0435